# Critical Filaments and Superconductivity in Quasiperiodic Twisted Bilayer Graphene

### v1

@@ Abstract
Moiré materials can exhibit electronic topological features yet are inherently quasiperiodic. Nonetheless, the localizing tendency of quasiperiodicity can be prevented by topology. We consider a quasiperiodic variant of the chiral Bistritzer-MacDonald model for twisted bilayer graphene with two incommensurate moiré potentials. We observe “filaments” linking magic angles with enhanced density of states and fractal wave functions that evade localization; states away from the filaments mimic fractal surface states of dirty topological superconductors. We demonstrate that topological quasiperiodicity can broadly enhance superconductivity without magic-angle fine-tuning.
@@
Amplifying electronic interactions by band-velocity engineering drives much of the research in moiré materials, exemplified by twisted bilayer graphene (TBLG) [1]. Adjusting the twist angle between layers tunes the carrier velocity in the moiré bands [2, 3], leading to (nearly) flat bands at certain magic angles [4, 5]. Band flattening strongly enhances interaction effects, and led to the observation of many correlated phases including superconductivity [1]. An additional feature of TBLG is the topological character of the effective bands [7, 8, 6, 9], which is enabled by the coupling of same-chirality Dirac electrons from the two layers. The confluence of strong interaction effects and topology at the magic angles facilitates the rich panorama of phenomena observed in TBLG, but also makes the experiments and the theoretical understanding thereof sensitive to fine details.
The quasiperiodic character of the microscopic structure leads to incommensurate tunneling with respect to the underlying lattices. At normal twist angles in TBLG showing correlated behavior, these effects are negligible. In contrast, incommensuration becomes more pronounced in trilayer systems where two moiré patterns can compete; superconductivity was very recently observed in such a quasiperiodic crystal [10]. Quasiperiodicity can even occur with respect to the often benign, unaligned subtrate: if twisted bilayer graphene is brought into alignment with the boron nitride substrate, one can see the emergence of quasicrystalline structure in scanning tunneling microscopy experiments [11].
@@
In topological materials, topology competes with the localizing nature of disorder in low-dimensions, protecting boundary and bulk states in nontrivial bands. Quasiperiodicity can also drive localization [12, 13]. The interplay between weak localization and interactions can sometimes produce surprising effects. For example, the average Cooper pairing amplitude can be enhanced near a superconductor-insulator transition if long-range Coulomb interactions are externally screened. The mechanism involves the Russian-doll-like nesting (“Chalker scaling”) of multifractal single-particle wave functions near the localization transition [14, 15], which boosts interaction matrix elements. A similar enhancement in self-consistent BCS mean-field theory was observed in quasiperiodic models [16, 17, 18].
@@
This raises an important question: can the incommensurate effects in moiré materials change the single-particle electronic wave functions in such a way that correlated behavior, such as superconductivity, can be enhanced? Here we numerically study a quasiperiodic variant of the chiral Bistritizer-MacDonald (BM) model for TBLG. One of the central results of this Letter is that incommensuration effects can robustly amplify superconducting pairing without the need to fine-tune to magic angles.
Previous work on incommensuration in Dirac and moiré materials focused on tight-binding models where the lattice provided incommensuration [19, 20, 21]; these works found single-particle, critical phases associated with “magic-angles” providing the first hint of multifractality. In contrast, the model considered here captures long wavelength physics where the moiré pattern (and the resulting super-moiré pattern) governs the low energy processes and topology emerges. We therefore expect our observations here are relevant for twisted stackings past the bilayer setup.
@@
The original BM model for twisted bilayer graphene [2, 4] possesses an emergent particle-hole symmetry [6], and it falls into class C in the $10$-fold classification of topological matter [22] and 2D Dirac Hamiltonians [23]. If AA tunneling vanishes, an additional chiral symmetry emerges [5]. This chiral BM model takes the same form as the 2D surface theory of a bulk class CI topological superconductor (TSC) [24]; the version with only particle-hole symmetry corresponds to the same topological surface, subject to orbital time-reversal symmetry breaking.
@@
It has long been known that quenched disorder which respects the chiral symmetry leads to remarkable results for class-CI topological Dirac surface states, including a power-law vanishing density of states (DOS) $\nu(E)\sim\left|E\right|^{1/7}$ [25] and critically Anderson delocalized low-energy wave functions with universal multifractal statistics [26]. Recent results have demonstrated that a different kind of universal wave-function criticality can extend over a wide range of the energy spectrum for these special dirty Dirac models (“spectrum-wide quantum criticality,” SWQC [27, 28]), although this can be disrupted at the 2D boundary of bulk topological systems by lattice-scale effects [29].
In this Letter, we explore the consequences of quasiperiodicity, revealing DOS enhancements and robust wave-function fractality, and demonstrate its connection to the disordered case. Using self-consistent BCS calculations [18], we show that the quasiperiodicity-induced fractality enhances superconducting pairing [14, 15] without fine-tuning to magic angles.
[…]
@@
Here $\tau^{\pm}=(\tau_{x}\pm i\tau_{y})/2$ (and similarly for $\sigma^{\pm}$) with $\tau_{j}$ as the Pauli matrices that link the layers. The potentials $U_{0}$ and $U_{1}$ take the form [4]
$U_{0}(\mathbf{r})$ $=e^{-i\mathbf{q}_{1}\cdot\mathbf{r}}+e^{-i\mathbf{q}_{2}\cdot\mathbf{r}}+e^{-i\mathbf{q}_{3}\cdot\mathbf{r}}\,$ (3)
$U_{1}(\mathbf{r})$ $=e^{-i\mathbf{q}_{1}\cdot\mathbf{r}}+e^{-i\frac{2\pi}{3}}e^{-i\mathbf{q}_{2}\cdot\mathbf{r}}+e^{i\frac{2\pi}{3}}e^{-i\mathbf{q}_{3}\cdot\mathbf{r}}\,.$ (4)
@@
Here $\mathbf{q}_{1}=k_{\theta}(0,1)$ and $\mathbf{q}_{2/3}=k_{\theta}(\mp\sqrt{3}/2,-1/2)$ are the reciprocal lattice vectors of the usual moiré potential. The momentum $k_{\theta}=2K\sin\theta/2$ defines the moiré scale (with $K$ as the distance from the K point to the $\Gamma$ point in the original graphene lattice). The second moiré potential $V_{\mathbf{q}^{\prime}}(\mathbf{r})$ takes the same form but with different reciprocal vectors $\{\mathbf{q}^{\prime}_{i}\}$ and potential strength $A_{2}$. When the two sets $\{\mathbf{q}_{i}\}$ and $\{\mathbf{q}^{\prime}_{i}\}$ are mutually incommensurate, we have a quasiperiodic model. In this Letter, we focus on the case $\mathbf{q}_{i}=\beta\,\mathbf{q}_{i}^{\prime}$ with $\beta$ the inverse golden ratio $(\sqrt{5}-1)/2$; we have verified that similar results obtain for other (e.g., silver) ratios [30].
The qBM model is realized in momentum space using a momentum cutoff $\Lambda=N_{\Lambda}k_{\theta}/2$. Focusing on $V_{\mathbf{q}}(\mathbf{r})$, the moiré Brillouin zone (mBZ) has linear size $k_{\theta}$, and we keep $2N_{\Lambda}^{2}$ moiré bands. Without a second potential, a momentum space lattice of linear size $N_{\Lambda}$ is sufficient; however, $\mathbf{q}^{\prime}_{i}$ allows hoppings within the mBZ. Since $|\mathbf{q}_{i}^{\prime}|=k_{\theta}/\beta$, we controllably reach larger system sizes with rational approximations of $\beta$, i.e. $\beta_{n}=F_{n}/F_{n+1}$ for Fibonacci numbers $F_{n}$. In this case, the denominator $\frac{|\mathbf{q}_{i}^{\prime}|}{k_{\theta}}=\frac{F_{n+1}}{F_{n}}$ determines the linear size of momentum space points required within one mBZ. Therefore, the full momentum space lattice has linear size $N_{\Lambda}F_{n}$ and spacing $\Delta k=k_{\theta}/F_{n}$; in the incommensurate limit ($n\rightarrow\infty$), all momentum can be accessed. In the following, we work in units such that $v_{F}=1$ and $k_{\theta}=\left|\mathbf{q}_{i}\right|=1$. Our setup is illustrated in Fig. 1.
@@
Figure 2: The zero-energy density of states $\nu(0)$ for the qBM model in chiral limit. Here $A_{1}$ and $A_{2}$ are the strengths of the two sets of moiré potentials. (a): The map of $\nu(0)$ for the two moiré potentials. The zero-energy DOS has large peaks at the magic angles of the $A_{1}$ and $A_{2}$ potentials [indicated by red stars in (a)]. The magic angles of $A_{1}$ and $A_{2}$ are connected by several filaments, where $\nu(0)$ peaks. (b), (c) and (d): $\nu(0)$ along the cuts at $A_{1}=0.4$, $1$ and $1.5$, respectively. The peaks correspond to the filaments and $\nu(0)$ remains nonzero between these. This is our first hint that quasi-periodicity has driven a single-particle phase transition away from the semimetal in these regions. The data are obtained with KPM for a momentum lattice with $N_{\Lambda}=17$ and $M=89$. We use Chebyshev expansion order $N_{\scriptscriptstyle\mathcal{C}}=2^{11}$ for the DOS map (a) and $N_{\scriptscriptstyle\mathcal{C}}=2^{11}$, $2^{12}$ and $2^{13}$ in (b-d).
@@
Density of states and fractal wave functions.—Dirac cones have a vanishing DOS $\nu(E)\sim|E|/v^{2}$ with velocity $v$; at the magic angle $v\rightarrow 0$ and $\nu(0)$ diverges. In the quasiperiodic case, at small $A_{1,2}$ the semimetal is stable and eigenstates remain plane-wave-like [36]. However, when $\nu(0)>0$ (or vanishes with a small power-law), multifractality–typically associated to an Anderson metal-insulator transition [37]–can occur. To reveal these (potential) single-particle phases, Fig. 2 shows $\nu(0)$ for the qBM model evaluated with the kernel polynomial method (KPM) [38]. Fig. 2(a) shows the $\nu(0)$ landscape while $A_{1,2}$ vary. There are large lakes with $\nu(0)>0$ separated by filaments where the DOS is strongly enhanced. The filaments connect the magic angles (red stars) of the monochromatic potentials; both filaments and lakes host multifractal wave functions (see below). Fig. 2(b-d) show particular cuts. In the regime close to the $A_{1}$ or $A_{2}$ axes away from the magic angles, the system remains semimetallic and $\nu(0)\rightarrow 0$. By contrast, $\nu(0)$ peaks at the filaments and either remains nonzero or vanishes with a nontrivial power-law between the filaments [30]. The system exhibits semi-metallic behavior for small $A_{1}$ or $A_{2}$ prior to the monochromatic magic angles. In the regions of comparable $A_{1}$ and $A_{2}$, between the filaments, $\nu(0)$ is generally nonzero; in fact, as we show below, the system exhibits SWQC behavior in these regions associated to class-CI surface states [27].
@@
As illustrated in Fig. 1, both cutoff $\Lambda$ and number of moiré unit cells $F_{n}$ determine the momentum space lattice, which need to both be large; therefore, to determine wave functions, we employ the Chebyshev filtering and shaking method [39, 40]. This method obtains eigenfunctions in a small energy window, which could contain thousands of states near a narrow band [30]. We use the method to obtain eigenfunctions for the qBM model with $\sim 10^{6}$ lattice sites.
@@
Figure 4: The second multifractal dimension $\tau_{2}$ versus energy and the wave functions at the one of filaments and in the lakes between the filaments. (a): $\tau_{2}$-$E$ at the filament with $A_{1}=0.4$ and $A_{2}=0.36$; (b): same as (a) but in the lakes between filaments with $A_{1}=0.4$ and $A_{2}=0.6$. (c)-(e): the wave functions at the filament for states at $E=0$, $0.2$ and $0.425$; (f)-(h): the wave functions in the lake for states at $E=0$, $0.2$ and $0.5$. The shadows in (a) and (b) show the energy gaps in the spectrum. The navy dashed and orange dot-dashed line in (a) and (b) show the reference value of $\tau_{2}$ for the WZW and SQHT states, respectively. Here we $N_{\Lambda}=11$, $M=55$ and the wave functions reside on a $605\times 605$ lattice in real space.
The statistics of single-particle wave function can be characterized by the inverse participation ratio (IPR),
$P_{q}=\sum_{i}\left|\psi_{i}\right|^{2q}\,,$ (5)
@@
Here $\psi_{i}$ is the (normalized) wave function in real space, obtained by Fourier transform of the momentum-space wave function. For extended wave functions, $P_{q}\sim L^{-d(q-1)}$, with $L$ the linear dimension of the system; for localized wave functions, $P_{q}\sim L^{0}$. Further, if $P_{q}\sim L^{-\tau_{q}}$ and $0<\tau_{q}<d(q-1)$, that wave function is critical. In the thermodynamic limit ($L\to\infty$), the multifractal dimension can be evaluated via
$\tau_{q}=-\log{P_{q}}/\log{L}\,.$ (6)
The $\tau_{q}$ can also be obtained with the box averaged IPR, $P_{q}^{\scriptscriptstyle\mathsf{box}}=\sum_{i_{b}}\left(\sum_{\text{box}}\left|\psi_{i}\right|^{2}\right)^{q}\propto b^{\tau_{q}}\,.$ with $b$ the linear size of the box. We use both methods to calculate the $\tau_{q}$ and they are consistent with each other for the appropriate choice of box sizes.
@@
The second multifractal dimension $\tau_{2}$ is shown in Fig. 3 for the qBM model along the cuts $A_{1}=0.4$ and $1$. Corresponding $\tau_{3}$ plots appear in [30]. Note that a value of $\tau_{2}\approx 2$ indicates ballistic wavefunctions, and Fig. 3(a) shows states like that for small $A_{2}$ and $A_{1}=0.4$. We observe a correlation between $\nu(0)$ and multifractality; at the first filament (with increasing $A_{2}$), $\tau_{2}$ decreases, fluctuates for stronger $A_{2}$, and dips near the additional filaments. The value of $\tau_{2}$ near the filaments in Fig. 3(a,c) indicates that the wave functions there are strongly multifractal, see also Fig. 4(c,d,g). Note that Fig. 3(c) follows the pattern for magic-angle semimetals as outlined in [20, 21]. In other regimes, such as between the filaments, the wave functions are also multifractal with a different value of $\tau_{2}$. In the nonperturbative regime, e.g., Fig. 3(b) along $A_{1}=1$, $\tau_{2}$ fluctuates around $3/2$ and is only partially correlated with $\nu(0)$. However, at this point $\tau_{2}$ appears to have saturated to the known value $\tau_{2}=3/2$ for critical states in class-CI Dirac models, where the emergent Wess-Zumino-Witten (WZW) term enables the zero-energy state to escape Anderson localization [26, 37, 23]. [Fig. 3(d) also illustrates this saturation.]
@@
Finite-energy surface-state wave functions of a dirty TSC can also evade localization via SWQC, with critical statistics tied to a quantum-Hall plateau transition [27, 41, 28, 29]. For the class-CI qBM model, we anticipate finite-energy states with SWQC associated to the spin quantum-Hall transition (SQHT) [27, 28]. To compare with finite energies in the qBM model, we take two separate values, on the filament ($A_{1}=0.4$ and $A_{2}=0.36$) and between the filaments ($A_{1}=0.4$ and $A_{2}=0.6$) [see Fig 4(a,b)], and examine a range of energies, extracting $\tau_{2}$. On the filament [Fig 4(a)], the low-energy states are strongly multifractal (with $\tau_{2}\approx 1$) with a few isolated peaks [energies $E=0,0.2$ in both Figs. 4(c,d)]; the higher energies oscillate between SQHT states [$\tau_{2}\approx 7/4$ and less strongly peaked, see Fig. 4(e)] and strongly multifractal states. The filaments, originating from the flat bands at the magic-angles, have hybridized those states strongly into these multifractal states. In between the filaments [Fig 3(b)], the low-energy wave functions are less multifractal than on the filament with WZW-like states [$\tau_{2}\approx 3/2$ and $\tau_{3}\approx 5/2$, Fig. 4(g)] and the high-energy states that reside deep within moiré bands are SQHT states [$\tau_{2}\approx 7/4$ and $\tau_{3}\approx 13/4$, Fig. 4(h)]. Near the energy gaps, the wave functions return to strongly multifractal or WZW-like states.
Figure 5: The superconductivity order parameter $\Delta$, zero energy DOS $\nu(0)$, distribution of the local order parameter $\Delta_{i}$ and superfluid stiffness for the qBM-BCS model. (a): $\Delta$ and $\nu(E=0)$ along the cut $A_{1}=0.4$; (b): same as (a) along the cut $A_{1}=1$; (c): the distribution of local order parameter $\Delta_{i}$ along $A_{1}=0.4$; (d) same as (c) along the cut $A_{1}=1$; (e): superfluid stiffness (red) along $A_{1}=0.4$ (clean limit value $2\Delta/\pi$ indicated by black dashed curve); (f): same as (e) along $A_{1}=1$.
Superconductivity.—The Bogoliubov-de Gennes Hamiltonian for the qBM model with pseudospin, layer, and spin-singlet $s$-wave pairing is given by
@@
Here $\chi$ is the Nambu spinor spanning sublattice ($\sigma$), layer ($\tau$), and particle-hole spaces ($\mu$) and

### v2

@@ Abstract
Multilayer moiré materials can exhibit topological electronic features yet are inherently quasiperiodic—leading to wave function interference whose Anderson-localizing tendency can be mitigated by topology. We consider a quasiperiodic variant of the chiral Bistritzer-MacDonald model for twisted bilayer graphene with two incommensurate moiré potentials that serves as a toy model for twisted trilayer. We observe “filaments” linking magic angles with enhanced density of states and fractal wave functions that evade localization; states away from the filaments mimic fractal surface states of dirty topological superconductors. We demonstrate that topological quasiperiodicity can broadly enhance superconductivity without magic-angle fine-tuning.
@@
Amplifying electronic interactions and uncovering topological phases drives much of the research in moiré materials [1]. Correlated phases including superconductivity were observed in twisted bilayer graphene (TBLG) [1], attributed to enhanced interactions due to flattening of the moiré bands near magic angles [2, 3, 4, 5]. Concurrently, topology is enabled in TBLG by the coupling of same-chirality Dirac electrons between the layers [7, 8, 6, 9].
Generic twist angles in moiré materials lead to incommensurate (quasiperiodic) tunneling, but this has negligible effects in TBLG. By contrast, quasiperiodicity is ubiquitous when multiple moiré patterns compete, as occurs in systems such as twisted trilayer graphene (TTLG), where superconductivity was recently observed [10, 11]. Previous work on quasiperiodic Dirac and moiré materials focused on tight-binding models with lattice incommensuration [12, 13, 14], and provided glimpses of a kind of wave-function criticality linked with “magic angles.” Wave-function criticality (multifractality) is typically associated to disorder-driven Anderson localization transitions [15]. Quasiperiodicity induces interference that can also drive Anderson localization, although the most-studied cases in one dimension (1D) show features different from random 1D systems [16, 17]. At the same time the tendency towards localization can be mitigated by topology, which protects boundary and bulk states in nontrivial bands.
@@
The interplay between multifractality and interactions can produce surprising effects. For example, the average Cooper pairing amplitude can be enhanced near a superconductor-insulator transition if Coulomb interactions are externally screened. The mechanism involves the Russian-doll-like nesting (“Chalker scaling”) of multifractal states, and can occur near the localization transition [18, 19] and in 1D quasiperiodic models [20, 21, 22].
Figure 1: (a): Pairing amplitude $\Delta$ and density of states (DOS) $\nu(E)$ for the chiral Bistritzer-MacDonald (BM) model; the enhancement of superconductivity occurs near the magic-angle ($A_{1}\approx 0.586$). (b): The same for the quasiperiodic BM model; superconductivity is broadly enhanced by wave-function fractality without magic-angle fine-tuning. The parameters $A_{1,2}$ are the strengths of the two incommensurate BM potentials, see Eqs. (1) and (2). All results are obtained from self-consistent BCS numerics using the kernel polynomial method [23]; $U$ is the interaction strength measured relative to the kinetic energy. The inset shows the $U$-dependence of $\Delta$ at $A_{1}=1$ and $A_{2}=0.6$. Here the combination of Chalker scaling and topology [24, 25, 26] leads to BCS-like $\Delta\sim e^{-c/U}$ (see text), instead of the threshold behavior of a 2D Dirac semimetal. The latter occurs away from the magic angle in (a).
@@
This raises a question: Can quasiperiodic and topological effects in multilayer moiré materials lead to criticality that enhances correlated behavior, such as superconductivity? We address this by studying a quasiperiodic variant of the chiral Bistritizer-MacDonald (BM) model for TBLG. Our model incorporates two incommensurate BM potentials as occurs in TTLG, and serves as a toy version of the latter with two layers instead of three. As we show in this Letter, quasiperiodicity with topology can lead to criticality and amplify pairing without the need to fine-tune to magic angles, see Fig. 1.
@@
The BM model for TBLG [2, 4] possesses an emergent particle-hole symmetry [6], falling into class C in the $10$-fold classification of topological matter [27] and 2D Dirac Hamiltonians [28]. If AA tunneling vanishes, a chiral symmetry emerges [5]. This chiral BM model takes the same form as the 2D surface theory of a bulk class CI topological superconductor (TSC) [29], while the BM model corresponds to the same surface, subject to orbital time-reversal symmetry breaking.
@@
Quenched disorder, which respects the chiral symmetry, leads to remarkable results for class-CI TSC surface states, including a power-law vanishing density of states (DOS) $\nu(E)\sim\left|E\right|^{1/7}$ [30] and critical low-energy wave functions with universal multifractal statistics [31]. Recent results have demonstrated that a more radical kind of universal wave-function criticality can extend over a wide range of the energy spectrum for these surfaces (“spectrum-wide quantum criticality,” SWQC [24, 25, 26]).
In this Letter, we show that topological quasiperiodicity leads to DOS enhancements and wave-function fractality. We establish that quasiperiodic multilayer graphene effectively simulates disordered TSC surface states. The resulting SWQC and absence of Anderson localization are exotic TSC features [24, 25] that are synthetically realizable in multilayer moiré materials. Our results are surprising because they precisely link the criticality of dirty surface states of bulk topological phases to quasiperiodic layered 2D materials. Since the bulk is “missing,” one might say that quasiperiodic twisted multilayers holographically realize bulk topological superconductor classes.
We study the consequences of topologically-protected quasiperiodicity for superconductivity using self-consistent BCS calculations [22], finding enhanced Cooper pairing [18, 19] without fine-tuning to magic angles (Fig. 1). For simplicity, we consider intravalley pairing. Our calculations incorporate additional aspects beyond those of the normal state such as Altshuler-Aronov corrections, which can also cause Anderson localization; such corrections shift the Bogoliubov-de Gennes eigenstate localization transition in a 1D quasiperiodic model [22]. We demonstrate that this effect is absent here due to topological protection [32], and we find that the superfluid stiffness precisely encodes the topological character of the normal state (another instance of the “holography”). Our approach can be applied to the more complicated case of TTLG [33].
@@
Here $\tau^{\pm}=(\tau_{x}\pm i\tau_{y})/2$ (and similarly for $\sigma^{\pm}$) with $\tau_{j}$ as the Pauli matrices that link the layers. The potentials $U_{0}$ and $U_{1}$ take the forms [4] $U_{0}(\mathbf{r})=e^{-i\mathbf{q}_{1}\cdot\mathbf{r}}+e^{-i\mathbf{q}_{2}\cdot\mathbf{r}}+e^{-i\mathbf{q}_{3}\cdot\mathbf{r}}$ and $U_{1}(\mathbf{r})=e^{-i\mathbf{q}_{1}\cdot\mathbf{r}}+e^{-i\frac{2\pi}{3}}e^{-i\mathbf{q}_{2}\cdot\mathbf{r}}+e^{i\frac{2\pi}{3}}e^{-i\mathbf{q}_{3}\cdot\mathbf{r}}\,.$ Here $\mathbf{q}_{1}=k_{\theta}(1,0)$ and $\mathbf{q}_{2/3}=k_{\theta}(-1/2,\pm\sqrt{3}/2)$ are the reciprocal lattice vectors of the usual moiré potential. The momentum $k_{\theta}=2K\sin\theta/2$ defines the moiré scale (with $K$ the K-point wavevector modulus in graphene). The second moiré potential $V_{\mathbf{q}^{\prime}}(\mathbf{r})$ takes the same form but with different reciprocal vectors $\{\mathbf{q}^{\prime}_{i}\}$ and potential strength $A_{2}$. When the two sets $\{\mathbf{q}_{i}\}$ and $\{\mathbf{q}^{\prime}_{i}\}$ are mutually incommensurate, we have a quasiperiodic model. In this Letter, we focus on the case $\mathbf{q}_{i}=\beta\,\mathbf{q}_{i}^{\prime}$ with $\beta$ the inverse golden ratio $(\sqrt{5}-1)/2$; we have verified that similar results obtain for other (e.g., silver) ratios [23].
@@
Figure 2: The zero-energy density of states $\nu(0)$ for the qBM model in chiral limit. Here $A_{1}$ and $A_{2}$ are the strengths of the two sets of moiré potentials. (a): The map of $\nu(0)$ for the two moiré potentials. The zero-energy DOS has large peaks at the magic angles of the $A_{1}$ and $A_{2}$ potentials [indicated by red stars in (a)]. The magic angles of $A_{1}$ and $A_{2}$ are connected by several filaments, where $\nu(0)$ peaks. (b), (c) and (d): $\nu(0)$ along the cuts at $A_{1}=0.4$, $1$ and $1.5$, respectively. The peaks correspond to the filaments and $\nu(0)$ remains nonzero between these. This is our first hint that quasi-periodicity has driven a single-particle phase transition away from the semimetal in these regions. The data are obtained with KPM for a momentum lattice with $N_{\Lambda}=17$ and $F_{n}=89$. We use Chebyshev expansion order $N_{\scriptscriptstyle\mathcal{C}}=2^{11}$ for the DOS map (a) and $N_{\scriptscriptstyle\mathcal{C}}=2^{11}$, $2^{12}$ and $2^{13}$ in (b-d).
The qBM model is realized in momentum space using a momentum cutoff $\Lambda=N_{\Lambda}k_{\theta}/2$. Focusing on $V_{\mathbf{q}}(\mathbf{r})$, the moiré Brillouin zone (mBZ) has linear size $k_{\theta}$, and we keep $2N_{\Lambda}^{2}$ moiré bands. Without a second potential, a momentum space lattice of linear size $N_{\Lambda}$ is sufficient; however, $\mathbf{q}^{\prime}_{i}$ allows hoppings within the mBZ. Since $|\mathbf{q}_{i}^{\prime}|=k_{\theta}/\beta$, we controllably reach larger system sizes with rational approximations of $\beta$, i.e. $\beta_{n}=F_{n}/F_{n+1}$ for Fibonacci numbers $F_{n}$. In this case, the denominator $\frac{|\mathbf{q}_{i}^{\prime}|}{k_{\theta}}=\frac{F_{n+1}}{F_{n}}$ determines the linear size of momentum space points required within one mBZ. Therefore, the full momentum space lattice has linear size $N_{\Lambda}F_{n}$ and spacing $\Delta k=k_{\theta}/F_{n}$; in the incommensurate limit ($n\rightarrow\infty$), all momentum can be accessed. In the following, we work in units such that $v_{F}=1$ and $k_{\theta}=\left|\mathbf{q}_{i}\right|=1$; all energies are measured in units of $v_{F}\,k_{\theta}$. Our setup is further illustrated in Ref. [23].
@@
Density of states and fractal wave functions.—Dirac cones have a vanishing DOS $\nu(E)\sim|E|/v^{2}$ with velocity $v$; at the magic angle $v\rightarrow 0$ and $\nu(0)$ diverges. In the quasiperiodic case, at small $A_{1,2}$ the semimetal is stable and eigenstates remain plane-wave-like [34]. However, when $\nu(0)>0$ (or vanishes with a small power-law), multifractality–typically associated to an Anderson metal-insulator transition [15]–can occur. To reveal these (potential) single-particle phases, Fig. 2 shows $\nu(0)$ for the qBM model evaluated with the kernel polynomial method (KPM) [35]. Fig. 2(a) shows the $\nu(0)$ landscape while $A_{1,2}$ vary. There are large lakes with $\nu(0)>0$ separated by filaments where the DOS is strongly enhanced. The filaments connect the magic angles (red stars) of the monochromatic potentials; both filaments and lakes host multifractal wave functions (see below). Fig. 2(b-d) show particular cuts. In the regime close to the $A_{1}$ or $A_{2}$ axes away from the magic angles, the system remains semimetallic and $\nu(0)\rightarrow 0$. By contrast, $\nu(0)$ peaks at the filaments and either remains nonzero or vanishes with a nontrivial power-law between the filaments [23]. In the regions of comparable $A_{1}$ and $A_{2}$, between the filaments, $\nu(0)$ is generally nonzero; in fact, as we show below, the system exhibits SWQC behavior in these regions associated to class-CI surface states [24].
@@
Both cutoff $\Lambda$ and number of moiré unit cells $F_{n}$ determine the momentum space lattice [23], which need to both be large; therefore, to determine wave functions, we employ the Chebyshev filtering and shaking method [36, 37]. This method obtains eigenfunctions in a small energy window, which could contain thousands of states near a narrow band [23]. We use the method to obtain eigenfunctions for the qBM model with $\sim 10^{6}$ lattice sites.
@@
Figure 4: The second multifractal dimension $\tau_{2}$ versus energy and the wave functions at the one of filaments and in the lakes between the filaments. (a): $\tau_{2}$-$E$ at the filament with $A_{1}=0.4$ and $A_{2}=0.36$; (b): same as (a) but in the lakes between filaments with $A_{1}=0.4$ and $A_{2}=0.6$. (c)-(e): the wave functions at the filament for states at $E=0$, $0.2$ and $0.425$; (f)-(h): the wave functions in the lake for states at $E=0$, $0.2$ and $0.5$. The shadows in (a) and (b) show the energy gaps in the spectrum. The navy dashed and orange dot-dashed line in (a) and (b) show the reference value of $\tau_{2}$ for the WZW and SQHT states, respectively. Here we $N_{\Lambda}=11$, $F_{n}=55$ and the wave functions reside on a $605\times 605$ lattice in real space.
@@
The statistics of single-particle wave functions can be characterized by the inverse participation ratio (IPR) $P_{q}\equiv\sum_{i}\left|\psi_{i}\right|^{2q}\,.$ Here $\psi_{i}$ is the (normalized) wave function in real space, obtained by Fourier transform of the momentum-space wave function. For extended wave functions, $P_{q}\sim L^{-d(q-1)}$, with $L$ the linear dimension of the system; for localized wave functions, $P_{q}\sim L^{0}$. Further, if $P_{q}\sim L^{-\tau_{q}}$ and $0<\tau_{q}<d(q-1)$, that wave function is critical. In the thermodynamic limit ($L\to\infty$), the multifractal dimension can be evaluated via $\tau_{q}=-\log{P_{q}}/\log{L}\,.$ The $\tau_{q}$ can also be obtained with the box-averaged IPR, $P_{q}^{\scriptscriptstyle\mathsf{box}}=\sum_{i_{b}}\left(\sum_{\text{box}}\left|\psi_{i}\right|^{2}\right)^{q}\propto b^{\tau_{q}}\,,$ with $b$ the linear size of the box. We use both methods to calculate the $\tau_{q}$ and they are consistent for the appropriate box sizes.
@@
The second multifractal dimension $\tau_{2}$ for low-energy wave functions is shown in Fig. 3 for the qBM model along the cuts $A_{1}=0.4$ and $1$. Corresponding $\tau_{3}$ plots appear in [23]. The value $\tau_{2}\approx 2$ indicates ballistic wavefunctions, and Fig. 3(a) shows such states for small $A_{2}$ and $A_{1}=0.4$. We observe a correlation between $\nu(0)$ and multifractality; at the first filament (with increasing $A_{2}$), $\tau_{2}$ decreases, fluctuates for stronger $A_{2}$, and dips near the additional filaments. The value of $\tau_{2}$ near the filaments in Fig. 3(a,c) indicates that the wave functions there are strongly multifractal, see also Fig. 4(c,d,g). Note that Fig. 3(c) follows the pattern for magic-angle semimetals as outlined in [13, 14]. In other regimes, such as between the filaments, the wave functions are also multifractal with a different value of $\tau_{2}$. In the nonperturbative regime, e.g., Fig. 3(b) along $A_{1}=1$, $\tau_{2}$ fluctuates around $3/2$ and is only partially correlated with $\nu(0)$. However, at this point $\tau_{2}$ appears to have saturated to the known value $\tau_{2}=3/2$ for zero-energy Wess-Zumino-Witten (WZW) surface states of a disordered class-CI TSC [31, 15, 28]. [Fig. 3(d) also shows this saturation.]
@@
Finite-energy surface-state wave functions of a dirty TSC can also evade localization via SWQC, with critical statistics tied to a quantum-Hall plateau transition [24, 38, 25, 26]. For the class-CI qBM model, we anticipate finite-energy states with SWQC associated to the spin quantum-Hall transition (SQHT) [24, 25]. To compare with finite energies in the qBM model, we take two separate values, on the filament ($A_{1}=0.4$ and $A_{2}=0.36$) and between the filaments ($A_{1}=0.4$ and $A_{2}=0.6$) [see Fig. 4(a,b)], and examine a range of energies, extracting $\tau_{2}$. On the filament [Fig. 4(a)], the low-energy states are strongly multifractal (with $\tau_{2}\approx 1$) with a few isolated peaks [energies $E=0,0.2$, Figs. 4(c,d)]; the higher energies oscillate between SQHT states [$\tau_{2}\approx 7/4$ and less strongly peaked, see Fig. 4(e)] and strongly multifractal states. The filaments, originating from the flat bands at the magic-angles, have hybridized those states strongly into these multifractal states. In between the filaments [Fig. 4(b)], the low-energy wave functions are less multifractal than on the filament with WZW-like states [$\tau_{2}\approx 3/2$ and $\tau_{3}\approx 5/2$, Fig. 4(g)] and the high-energy states that reside deep within moiré bands are SQHT states [$\tau_{2}\approx 7/4$ and $\tau_{3}\approx 13/4$, Fig. 4(h)]. Near the energy gaps, the wave functions show stronger fractal features.